# LiveCodeBench: Holistic and Contamination Free Evaluation of Large Language Models for Code

### v1

@@ Abstract
Large Language Models (LLMs) applied to code-related applications have emerged as a prominent field, attracting significant interest from both academia and industry. However, as new and improved LLMs are developed, existing evaluation benchmarks (e.g., HumanEval, MBPP) are no longer sufficient for assessing their capabilities. In this work, we propose LiveCodeBench, a comprehensive and contamination-free evaluation of LLMs for code, which continuously collects new problems over time from contests across three competition platforms, namely LeetCode, AtCoder, and CodeForces. Notably, our benchmark also focuses on a broader range of code-related capabilities, such as self-repair, code execution, and test output prediction, beyond just code generation. Currently, LiveCodeBench hosts four hundred high-quality coding problems that were published between May 2023 and February 2024. We have evaluated base LLMs and instruction-tuned LLMs on LiveCodeBench. We present empirical findings on contamination, holistic performance comparisons, potential overfitting in existing benchmarks as well as individual model comparisons. We will release all prompts and model completions for further community analysis, along with a general toolkit for adding new scenarios and models.
@@ 1 Introduction
Live updates to prevent contamination. LLMs are trained on massive inscrutable corpora, and current benchmarks suffer from the risk of data contamination as they could be included in those training datasets. While previous works have attempted decontamination using both exact and fuzzy matches (Li et al., 2023b, d), it can be a non-trivial task (Team, 2024) and can be evaded using simple strategies like rephrasing (Yang et al., 2023). Here, to prevent the risk of problem contamination, we use live updates, that is evaluate models on new problems. Particularly, we collect problems from weekly contests on competition platforms and tag them with a release date. Next, for newer models, we only consider problems released after the model’s cutoff date to ensure that the model has not encountered the exact problem in the training dataset. In Figure 1, we find that the performance of the DeepSeek model starkly drops when evaluated on the LeetCode problems released after August . This indicates that DeepSeek is potentially trained on the older LeetCode problems.
@@
Holistic Evaluation. Current code evaluations primarily focus on natural language to code generation. However, programming is a multi-faceted task that requires a variety of capabilities beyond those measured by code generation. In LiveCodeBench, we evaluate code LLMs three additional scenarios, listed below.
@@
High-quality problems and tests. High-quality problems and tests are crucial for reliable evaluation of LLMs. However, prior works have revealed deficiencies in existing benchmarks. Liu et al. (2023a) identified insufficient tests and ambiguous problem descriptions in HumanEval. They released HumanEval+, a variant of the benchmark with more tests and sometimes saw up to an 8% drop in performance. Similarly, Austin et al. (2021) had to create a sanitized MBPP subset to disambiguate problem descriptions. In LiveCodeBench, we source the problems from reputable competition websites whose quality is already validated by the platform users. In addition, for every problem, we provide a large number of tests (over on average) for meaningful and robust evaluations.
@@
Balanced problem difficulty. Competition programming is challenging for even the best-performing LLMs, and most of the current SoTA models achieve close to zero performance on a majority of problems. As a result, they can be unsuitable for meaningful comparing today’s LLMs because the variance in performances is low. Furthermore, the averaging of evaluation scores across problems with different difficulty levels artificially minimizes the differences between models. In contrast, we use problem difficulty ratings (sourced from the competition websites) to ensure balanced problem difficulty distribution and allow granular model comparisons.
@@
Particularly, we have collected problems from contests across three competition platforms – LeetCode, AtCoder, and CodeForces occurring from May to the present (February ) and use them to construct the different LiveCodeBench scenarios.
@@
Contamination. We have observed a stark drop in the performance of DeepSeek on LeetCode problems released after Aug , highlighting likely contamination in older problems.
@@
Holistic Evaluation. Our evaluations reveal that model performances are correlated across tasks, but the relative differences do vary. For example, the gap between open and closed models further increases on tasks like self-repair or test output prediction. Similarly, certain models like Claude-3-Opus and Mistral-L perform considerably better on code execution and test output prediction compared to code generation with Claude-3-Opus surpassing GPT-4 on the test output prediction. This highlights the importance of a holistic evaluation.
@@
Among the open access base models, we find that DeepSeek-Base models are the strongest, followed by StarCoder2-Base and CodeLLaMa-Base. Their respective finetuned models, DS-Ins-33B and Phind-34B achieve strong performance.
Closed API models such as GPT, Claude, and Gemini generally outperform open models (Figure 2). The open models that cross the performance gap are DS-Ins-33B and Phind-34B, instruction-tuned variants of large base models (both have B parameters).
Existing benchmarks insufficiently highlight the gap between GPT-4 and other models. Particularly, smaller models achieve similar or often even better performance compared to GPT-4. In LiveCodeBench, GPT-4 (and GPT-4-Turbo) outperforms all other models (except Claude-3-Opus) by a large margin in all scenarios.
Concurrent Work. Huang et al. (2023) also evaluate LLMs in a time-segmented manner. However, they only focus on CodeForces problems while we combine problems across platforms and additionally propose a holistic evaluation across multiple code-related scenarios. Li et al. (2023c) propose a large dataset of competitive programming problems with additional generated tests but do not study contamination or tasks beyond generation. Liu et al. (2024) evaluate the code comprehension capabilities of LLMs using execution. Singhal et al. (2024) also propose evaluating LLMs on tasks beyond code generation, but they consider tasks that take into account the non-functional-correctness aspects of programming. Guo et al. (2024) also evaluate DeepSeek on LeetCode problems and mention the possibility of problem contamination.
@@ 3.1 Data Collection
Test collection. Tests are crucial for assessing the correctness of the generated outputs and are used in all four scenarios. We collect public-facing tests whenever possible and use them for the benchmark. Otherwise, following Liu et al. (2023b), we use a LLM (here GPT-4-Turbo) to generate tests for the problems. A key difference between our test generation approach and theirs is that instead of generating inputs directly using the LLM, we construct generators that sample inputs based on the problem specifications using few-shot prompting. Details and examples of such input generators can be found in Section A.2.
Problem difficulty. Competition programming has remained a challenge for LLMs, with GPT-4 achieving an average CodeForces rating (ELO) of 392, placing it in the bottom 5 percentile (OpenAI, 2023). This makes it difficult to compare LLMs, as the variation in performance across models is low. In LiveCodeBench, we collect problems of diverse difficulties as labeled in competition platforms, excluding problems that are rated above a certain threshold that are likely too difficult for even the best models. Further, we use these ratings to classify problems as Easy, Medium, and Hard for more granular model comparisons.
@@
| LCB (May-Feb) | 400 | 142 | 168 | 90 | 59.2 |
| LCB (Sep-Feb) | 238 | 85 | 98 | 55 | 56.9 |
[…]
| LeetCode | 181 | 62 | 90 | 29 | 96.3 |
| CodeForces | 9 | 4 | 2 | 3 | 44.3 |
[…]
| LCB-Medium | 168 | 0 | 168 | 0 | 66.8 |
| LCB-Hard | 90 | 0 | 0 | 90 | 57.8 |
@@ 3.2 Platform Specific Curation
LeetCode. We collect problems from all weekly and biweekly contests on LeetCode that have taken place since May’23. For each problem, we collect the problems, public tests, and user solutions. The platform also provides a difficulty label for each problem which we use to tag the problems as Easy, Medium, and Hard. Since LeetCode provides a starter code for each problem, we also collect it and provide it to the LLM in the STDIN format. Since the hidden tests are not directly available, we use our generator-based test input generation approach (Section A.2).
@@ 3.3 Scenario-specific benchmark construction
Code Execution. We draw inspiration from the benchmark creation procedure used in Gu et al. (2024). First, we collect a large pool of correct, human-submitted solutions from the LeetCode subset. However, many of these programs have multiple nested loops, complex numerical computations, and a large number of execution steps. Therefore, we apply compile-time and run-time filters to ensure samples are reasonable, and we double-check this with a manual inspection. More details on the filtering criteria and statistics of the dataset can be found in Appendix A.3. Our final dataset consists of samples from problems111Code execution scenario currently uses problems between May-November.
@@ 4 Experiment Setup
Models. We evaluate models across various sizes, ranging from B to B, including base models, instruction models, and both open and closed models. Our experiments include models from different classes, such as GPTs (GPT-3.5-turbo, GPT-4, GPT-4-Turbo), Claudes (Claude-Ins-1, Claude-2), Gemini-Pro among closed-access and DeepSeeks (DS-Ins-{1.3, 6.7, 33}B, DS-Base-{1.3, 6.7, 33}B), CodeLLaMas (CL-Ins-{7, 13, 34}B, CL-Base-{7, 13, 34}B), StarCoder2 (SC2-Base-{7,15}B) among open. Additionally, we also include fine-tuned models WizardCoder (WC-34B, WC-33B) and Phind-34B from CL-Base-34B, and MagiCoders (MC-6.7, 7B) from CL-Base-7B and DS-Base-6.7B. See Appendix C.1 for their details and cutoff or release dates.
Evaluation Metrics. We use the Pass@1 (Kulal et al., 2019; Chen et al., 2021) metric for our evaluations. Specifically, we generate candidate answers for each problem either using API or using vLLM (Kwon et al., 2023). We use nucleus sampling with temperature and top_p and calculate the fraction answers that are correct. For the code generation and self-repair scenarios, we use tests to verify the correctness of the programs. For the code execution scenario, we use an execution-based correctness metric between the generated output and the ground truth output. For the test output prediction scenario, we parse the generated response to extract the answer and use equivalence checks for grading as specified in Section 2.
@@ 4.1 Scenario-specific setup
Code Generation. For the instruction-tuned models, we use a zero-shot prompt and follow the approach of Hendrycks et al. (2021) by adding appropriate instructions to generate solutions in either functional or stdin format. For the base models, we use a constant one-shot example, with a separate example provided for problems that accept stdin input and for problems that accept functional output. Section C.2 shows the high-level zero-shot prompt used.
Self Repair. Similar to prior work Olausson et al. (2023), we use the programs generated during the code generation scenario along with the corresponding error feedback to build the zero-shot prompt for the self-repair scenario. The type of error feedback includes syntax errors, runtime errors, wrong answers, and time-limit errors, as applicable. Section C.3 provides the pseudo-code for computing the error feedback and the corresponding prompt.
Code Execution. We use few-shot prompts for the code execution scenario, both with and without chain-of-thought prompting (COT). Particularly, we use a 2-shot prompt without COT and a 1-shot prompt with COT with manually detailed steps. The prompts are detailed in Section C.4.
Test Output Prediction. We use a zero-shot prompt that queries the model to complete assertions, given the problem, function signature, and test input. We provide the prompt in Section C.5.
@@ 5.1 Avoiding Contamination
Contamination in DeepSeek. LiveCodeBench comprises problems released since May . However, DeepSeek models were released in Sep and might have already been trained on some of the problems in our benchmark. We can measure the performance of the model on the benchmark using problems released after the cutoff date, thereby estimating the performance of the model on previously unseen problems. Figure 1 shows the performance of DS-Ins-33B on LiveCodeBench code generation and test output prediction scenario on LeetCode problems released in different months from May and Jan . We notice a stark drop in performance after Aug. (right before its release date), which suggests that the earlier problems might indeed be contaminated. This trend is consistent across other LiveCodeBench scenarios like repair and code execution, as depicted in Figure 10. Concurrently, Guo et al. (2024) (Section 4.1, last paragraph) also acknowledge the possibility of LeetCode contamination, noting that “models achieved higher scores in the LeetCode Contest held in July and August”.
Interestingly, we find that this drop in performance primarily occurs for the LeetCode problems only and that the model performance is relatively smooth across the months for problems from other platforms. Figure 11 shows a relatively stable performance for all models on AtCoder problems released over different periods, with the possible exception of May and June.
Performances of other models. We study performance variations in other models released recently. Particularly, GPT-4-Turbo, Gemini-Pro, Mistral-L, and Claude-3s models were released in November , December , February , and March respectively. Note that GPT-4-Turbo (-preview variant) and Claude-3s have cutoff dates April and August respectively. Irrespective of the release or cutoff dates, we do not find any drastic performance variations across the months, as shown in Figure 12, particularly compared to the DeepSeek models. Interestingly, we find that even the DS-Base-33B model also suffers from contamination dropping from Pass@1 in May problems to Pass@1 in September LeetCode problems. This also suggests the likely inclusion of competition problems in the pretraining of the DeepSeek models.

### v2

@@ Abstract
Large Language Models (LLMs) applied to code-related applications have emerged as a prominent field, attracting significant interest from both academia and industry. However, as new and improved LLMs are developed, existing evaluation benchmarks (e.g., HumanEval, MBPP) are no longer sufficient for assessing their capabilities. In this work, we propose LiveCodeBench, a comprehensive and contamination-free evaluation of LLMs for code, which collects new problems over time from contests across three competition platforms, namely LeetCode, AtCoder, and CodeForces. Notably, our benchmark also focuses on a broader range of code-related capabilities, such as self-repair, code execution, and test output prediction, beyond just code generation. Currently, LiveCodeBench hosts over five hundred coding problems that were published between May 2023 and May 2024. We have evaluated base LLMs and instruction-tuned LLMs on LiveCodeBench. We present empirical findings on contamination, holistic performance comparisons, potential overfitting in existing benchmarks as well as individual model comparisons. We will release all prompts and model completions for further community analysis, along with a general toolkit for adding new scenarios and models.
@@ 1 Introduction
Live updates to prevent contamination. LLMs are trained on massive inscrutable corpora, and current benchmarks suffer from the risk of data contamination as they could be included in those training datasets. While previous works have attempted decontamination using both exact and fuzzy matches (Li et al., 2023b, d), it can be a non-trivial task (Team, 2024) and can be evaded using simple strategies like rephrasing (Yang et al., 2023). Here, to prevent the risk of problem contamination, we use live updates, that is evaluate models on new problems. Particularly, we collect problems from weekly contests on competition platforms and tag them with a release date. Next, for newer models, we only consider problems released after the model’s cutoff date to ensure that the model has not encountered the exact problem in the training dataset. In Figure 1, we find that the performance of the DeepSeek model starkly drops when evaluated on the LeetCode problems released after August . Similarly, GPT-4-O observes a drop in performance on LeetCode problems released since November 2023, its specified cutoff date. This indicates that these models are likely trained on the older LeetCode problems and time-segmented evaluations allow fair comparisions.
@@
Holistic Evaluation. Current code evaluations primarily focus on natural language to code generation. However, programming is a multi-faceted task that requires a variety of capabilities beyond those measured by code generation. In LiveCodeBench, we evaluate code LLMs on three additional scenarios, listed below.
@@
High-quality problems and tests. High-quality problems and tests are crucial for reliable evaluation of LLMs. However, prior works have revealed deficiencies in existing benchmarks. (Liu et al., 2023a) identified insufficient tests and ambiguous problem descriptions in HumanEval. They released HumanEval+, a variant of the benchmark with more tests and sometimes saw up to an 8% drop in performance. Similarly, (Austin et al., 2021) had to create a sanitized MBPP subset to disambiguate problem descriptions. In LiveCodeBench, we source the problems from reputable competition websites whose quality is already validated by the platform users. In addition, for every problem, we provide a good number of tests (about on average) for meaningful and robust evaluations while still finishing quickly.
@@
Balanced problem difficulty. Competition programming is challenging for even the best-performing LLMs, and most of the current SoTA models achieve close to zero performance on a majority of problems. As a result, they can be unsuitable for meaningful comparing today’s LLMs because the variance in performances is low. Furthermore, the averaging of evaluation scores across problems with different difficulty levels artificially minimizes the differences between models. Therefore, we use problem difficulty ratings (sourced from the competition websites) for filtering the harder problems and classifying problem difficulties to ensure balanced problem difficulty distribution and allow granular model comparisons.
@@
Particularly, we have collected problems from contests across three competition platforms – LeetCode, AtCoder, and CodeForces occurring from May to the present (May ) and use them to construct the different LiveCodeBench scenarios.
@@
Contamination. We observe a stark drop in the performance of DeepSeek, GPT-4-O, and Codestral on LeetCode problems released after Aug , Oct , and Jan (Figure 1). These results highlight likely contamination in older problems and time-segmented evaluations prove effective for performing fair comparisons.
@@
Holistic Evaluation. Our evaluations reveal that model performances are correlated across tasks, but the relative differences do vary. For example, in Figure 2, the gap between open and closed models further increases on tasks like self-repair or test output prediction. Similarly, Claude-3-Opus and Mistral-L perform considerably better on code execution and test output prediction compared to code generation with Claude-3-Opus surpassing GPT-4 on the test output prediction. This highlights the importance of a holistic evaluation.
@@
Among the open access base models, we find that L3-Base and DeepSeek-Base models are the strongest, followed by StarCoder2-Base and CodeLLaMa-Base.
Closed API models such as GPTs, Claude, and Gemini generally outperform open models (Figure 2). The open models that close the gap are L3-Ins-70B, Mixtral, and DS-Ins-33B are instruction-tuned variants of large base models (B parameters).
Existing benchmarks are insufficient at highlighting the gap between GPT-4 and other models. Particularly, smaller models achieve similar or often even better performance compared to GPT-4. In LiveCodeBench, GPT-4 (and GPT-4-Turbo) outperforms all other models (except Claude-3-Opus) by a large margin in all scenarios.
Concurrent Work. (Huang et al., 2023) also evaluate LLMs in a time-segmented manner. However, they only focus on CodeForces problems while we combine problems across platforms and additionally propose a holistic evaluation across multiple code-related scenarios. (Li et al., 2023c) propose a large dataset of competitive programming problems with additional generated tests but do not study contamination or tasks beyond generation. Liu et al. (2024) evaluate the code comprehension capabilities of LLMs using execution. (Singhal et al., 2024) also propose evaluating LLMs on tasks beyond code generation, but they consider tasks that take into account the non-functional-correctness aspects of programming. (Guo et al., 2024) also evaluate DeepSeek on LeetCode problems and mention the possibility of problem contamination.
@@ 3.1 Data Collection
Test collection. Tests are crucial for assessing the correctness of the generated outputs and are used in all four scenarios. We collect tests available on platform websites whenever possible and use them for the benchmark. Otherwise, following Liu et al. (2023b), we use a LLM (here GPT-4-Turbo) to generate tests for the problems. A key difference between our test generation approach is that instead of generating inputs directly using the LLM, we construct generators that sample inputs based on the problem specifications using in context learning. Details and examples of such input generators can be found in Section A.2. Finally, we collect a small fraction of failing tests from the platform for the more recent problems allowing more directed adversarial test collection.
Problem difficulty. Competition programming has remained a challenge for LLMs, with GPT-4 achieving an average CodeForces rating (ELO) of 392, placing it in the bottom 5 percentile (OpenAI, 2023). This makes it difficult to compare LLMs, as the variation in performance across models is low. In LiveCodeBench, we collect problems of diverse difficulties as labeled in competition platforms, excluding problems that are rated above a certain threshold that are likely too difficult for even the best models111From our early explorations, we find CodeForces problems being considerably more difficult than AtCoder and LeetCode problems and thus focus primarily on the latter platforms.. Further, we use these ratings to classify problems as Easy, Medium, and Hard for more granular model comparisons.
@@
| LCB (May-end) | 511 | 182 | 206 | 123 | 17.0 |
| LCB (Sep-end) | 349 | 125 | 136 | 88 | 18.0 |
| AtCoder | 267 | 99 | 91 | 77 | 15.6 |
| LeetCode | 235 | 79 | 113 | 43 | 19.0 |
| CodeForces | 9 | 4 | 2 | 3 | 11.1 |
| LCB-Easy | 182 | 182 | 0 | 0 | 16.1 |
| LCB-Medium | 206 | 0 | 206 | 0 | 17.4 |
| LCB-Hard | 123 | 0 | 0 | 123 | 18.0 |
@@ 3.2 Platform Specific Curation
LeetCode. We collect problems from all weekly and biweekly contests on LeetCode that have taken place after April’23. For each problem, we collect the problems, public tests, and user solutions. The platform also provides a difficulty label for each problem which we use to tag the problems as Easy, Medium, and Hard. Since LeetCode provides a starter code for each problem, we also collect it and provide it to the LLM in the STDIN format. Since the hidden tests are not directly available, we use our generator-based test input generation approach (Section A.2) and also collect the auto grader failing tests for some of the recent problems.
@@ 3.3 Scenario-specific benchmark construction
Code Execution. We draw inspiration from the benchmark creation procedure used in Gu et al. (2024). First, we collect a large pool of correct, human-submitted solutions from the LeetCode subset. However, many of these programs have multiple nested loops, complex numerical computations, and a large number of execution steps. Therefore, we apply compile-time and run-time filters to ensure samples are reasonable, and we double-check this with a manual inspection. More details on the filtering criteria and statistics of the dataset can be found in Appendix A.3. Our final dataset consists of samples from problems.
@@ 4 Experiment Setup
Models. We evaluate models across various sizes, ranging from B to B, including base models, instruction models, and both open and closed models. Our experiments include models from different classes, such as GPTs (GPT-3.5-turbo, GPT-4, GPT-4-Turbo,GPT-4-O), Claudes (Claude-Ins-1, Claude-2, Claude-3s), Geminis(Gemini-Pro, Gemini-Flash), Mistral among closed-access and LLaMa-3s(L3-Base-{7, 70}B, L3-Ins-{7, 70}B), DeepSeeks (DS-Base-{1.3, 6.7, 33}B, DS-Ins-{1.3, 6.7, 33}B), CodeLLaMas (CL-Ins-{7, 13, 34}B, CL-Base-{7, 13, 34}B), StarCoder2 (SC2-Base-{3,7,15}B), CodeQwen among open. Additionally, we also include fine-tuned models Phind-34B from CL-Base-34B, and MagiCoders (MC-{6.7, 7}B) from CL-Base-7B and DS-Base-6.7B. See Appendix C.1 for a complete list of models and estimated cutoff dates.
Evaluation Metrics. We use the Pass@1 (Kulal et al., 2019; Chen et al., 2021) metric for our evaluations. Specifically, we generate candidate answers for each problem either using API or using vLLM (Kwon et al., 2023). We use nucleus sampling with temperature and top_p and calculate the fraction of programs or answers that are correct. For the code generation and self-repair scenarios, we use tests to verify the correctness of the programs. For these scenarios, programs must pass all tests to be considered correct. For the code execution scenario, we use an execution-based correctness metric between the generated output and the ground truth output. For the test output prediction scenario, we parse the generated response to extract the answer and use equivalence checks for grading as specified in Section 2.
@@ 4.1 Scenario-specific setup
Code Generation. For the instruction-tuned models, we use a zero-shot prompt and follow the approach of Hendrycks et al. (2021) by adding appropriate instructions to generate solutions in either functional or stdin format. For the base models, we use a constant one-shot example, with a separate example provided for problems that accept stdin input and for problems that accept functional output. Section LABEL:appendix:codegen-prompts shows the high-level zero-shot prompt used.
Self Repair. Similar to prior work Olausson et al. (2023), we use the programs generated during the code generation scenario along with the corresponding error feedback to build the zero-shot prompt for the self-repair scenario. The type of error feedback includes syntax errors, runtime errors, wrong answers, and time-limit errors, as applicable. Section LABEL:appendix:repair-prompts provides the pseudo-code for computing the error feedback and the corresponding prompt.
Code Execution. We use few-shot prompts for the code execution scenario, both with and without chain-of-thought prompting (COT). Particularly, we use a 2-shot prompt without COT and a 1-shot prompt with COT with manually detailed steps. The prompts are detailed in Section LABEL:appendix:code-execution-prompts.
Test Output Prediction. We use a zero-shot prompt that queries the model to complete assertions, given the problem, function signature, and test input. We provide the prompt in Section LABEL:appendix:test-pred-prompts.
@@ 5.1 Avoiding Contamination
Contamination in DeepSeek and GPT-4-O. LiveCodeBench comprises problems released since May . However, DeepSeek models were released in Sep and might have already been trained on some of the problems in our benchmark. Similarly, OpenAI notes GPT-4-O cutoff date in November. We can measure the performance of the models on the benchmark using problems released after the cutoff date, thereby estimating the performance of the model on previously unseen problems. Figure 1 shows the performance of these models on LiveCodeBench code generation and test output prediction scenario on LeetCode problems released in different months from May and Feb . We notice a stark drop in the performance of DS-Ins-33B model after Aug. (right before its release date), which suggests that the earlier problems might indeed be contaminated. This trend is consistent across other LiveCodeBench scenarios like repair and code execution, as depicted in Figure LABEL:fig:contamination_all_tasks_line. Concurrently, Guo et al. (2024) (Section 4.1, last paragraph) also acknowledge the possibility of LeetCode contamination, noting that “models achieved higher scores in the LeetCode Contest held in July and August. Similarly, performance of the GPT-4-O model drops on problems released since November (its official cutoff date).
Interestingly, we find that this drop in performance primarily occurs for the LeetCode problems only and that the model performance is relatively smooth across the months for problems from other platforms. Figure LABEL:fig:atcoder_contamination shows a relatively stable performance for all models on AtCoder problems released over different periods, with the possible exception of May and June.
Performances of other models. We study performance variations in other models released recently. Particularly, GPT-4-Turbo, Gemini-Pro, Mistral-L, and Claude-3s models were released in November , December , February , and March respectively. Note that GPT-4-Turbo (-preview variant) and Claude-3s have cutoff dates April and August respectively. Irrespective of the release or cutoff dates, we do not find any drastic performance variations across the months, as shown in Figure LABEL:app:fig:contamination_all_models, particularly compared to the DeepSeek models. Interestingly, we find that even the DS-Base-33B model also suffers from contamination dropping from Pass@1 in May problems to Pass@1 in September LeetCode problems. This also suggests the likely inclusion of competition problems in the pretraining of the DeepSeek models, thereby affecting all instruction models trained from it. Finally, Codestral achieves Pass@1 on problems released between May’23 and Jan’24 and Pass@1 on problems since Feb’24.